Keziryahward014:

What is the solution for the equation 4x - 10 =2x

XxXNessalulbaddieXxX:

haiiii i can help

XxXNessalulbaddieXxX:

do u knw how to start this problem and the few steps u do how to do i can walk u through it

carmelle:

hi there! Lets do this step-by-step :) Given: 4x - 10 =2x are you familiar with solving for x?

xxDeppressionxx:

Yup I just seen that.

xxDeppressionxx:

4x−10=2x −10=2x−4x −10=−2x -10/-2=x 10/2=x 5=x X=5